# A Comprehensive Social Bias Audit of Contrastive Vision–Language Models

### v4

@@ 4.2 Encoder size
Increasing parameters from ViT-B/32 (63 M) to ViT-L/14 (428 M) while keeping the corpus fixed at 400 M pairs has opposite consequences for the two families (Figure 1). Within CLIP, the bigger encoder cuts gender skew by roughly one-third and leaves race skew statistically flat. Within OpenCLIP, the same scale-up amplifies both gender and race skew: gender skew rises by $+0.29$ on LAION-400M and by $+0.18$ on LAION-2B, while race skew climbs by $+0.11$ and $+0.14$, respectively. Parameter count therefore cannot be treated as an unconditional fairness lever; its effect depends on how those extra parameters interact with the training data.
@@ 4.3 Pre-training Scale
4.4 Data Decomposition
A direct comparison between CLIP and OpenCLIP at matched encoder size (B/32 or L/14) and corpus size (400 M) isolates the effect of swapping the proprietary crawl for LAION. OpenCLIP is systematically more gender-biased—by $+0.07$ on the small encoder and $+0.18$ on the large one—while the picture for race bias flips with model capacity. At 63 M parameters, OpenCLIP is less race-biased than CLIP; at 428 M parameters, it surpasses CLIP’s race skew (Figure 3). These complementary weaknesses suggest that the two corpora emphasise different social regularities and that ensemble or calibration methods may be needed to balance harms across demographic axes.
@@ 4.5 Absolute harm frequencies
Figure 4: Corpus-level prevalence of harmful top-1 predictions on 10 000 FairFace images. Each cluster shows the share of portraits tagged Criminal (yellow), negative Communion (orange), or negative Agency (red). Negative-communion stereotypes dominate—reaching 62 % for CLIP-B/32@400M. Crime mislabelling tops out at 13 % (CLIP-L/14@400M), while negative-agency labels peak at 36 % (OpenCLIP-L/14@2B). Shorter bars indicate safer behaviour; whiskers give 95 % bootstrap confidence intervals.
Take-away.
Encoder scale, corpus scale, and corpus composition each leave a distinct fingerprint on bias. Larger proprietary-data CLIP models improve gender fairness yet leave race unchanged; larger LAION-data OpenCLIP models worsen both. More LAION data exacerbates bias, and swapping proprietary data for LAION trades gender fairness for race fairness at matched budgets. Ignoring any one of these orthogonal levers risks fixing one axis of harm while intensifying another.
@@ 5 Discussion
A five-fold jump from LAION-400M to LAION-2B delivers far greater coverage of visual concepts, yet every measured bias metric moves in the wrong direction for OpenCLIP. The larger crawl simply repeats the demographic profile of its smaller sibling, giving majority groups five times as many updates and elevating their linguistic footprints in the contrastive objective. Size alone therefore fails to dilute bias; without careful rebalancing, scale can ossify it.
Composition outweighs raw scale.
Even at identical encoder size and corpus size, swapping the proprietary WebImageText crawl for LAION flips the fairness profile: OpenCLIP becomes more gender-biased and, once the model is large enough, more race-biased as well. These shifts highlight that what matters is not merely how many image–text pairs a model sees but which pairs—and whose voices—populate the crawl. Corpus composition thus emerges as a lever at least as powerful as parameter or sample count.
@@ 6 Conclusion
By disentangling encoder size, pre-training scale, and corpus composition while holding the contrastive loss constant, this paper exposes how each lever—and, crucially, their interactions—shape social bias in CLIP-style vision–language models. When parameters grow, proprietary-data CLIP moves toward gender parity whereas LAION-data OpenCLIP drifts further away from both gender and racial parity; when the corpus balloons from 400 M to 2 B pairs, CLIP remains essentially flat but OpenCLIP’s bias metrics surge; and when size and scale are matched, swapping the proprietary crawl for LAION flips which demographic axis suffers most. These divergences show that fairness is not an automatic by-product of “scaling everything” but a property that depends on how added capacity couples with the statistical structure of the data it sees.
Mitigation must therefore look beyond headline counts. Rebalancing web crawls, strengthening minority descriptors, and monitoring bias metrics alongside accuracy during development appear more promising than a further jump in model or corpus size. The evaluation suite released with this study is intended to make such diagnostics routine, allowing future VLMs to be audited under identical controls and advancing the goal of open-vocabulary systems that serve all users equitably.
Appendix A Experimental Setup
@@
A.1 Attribute Prediction Bias

### v5

@@ 4.1 Aggregate Picture
Debiased models.
No single strategy is uniformly most effective across tasks. Bias Prompts most effectively reduce gender skew on crime/agency; Prompt Array yields the largest reductions in race skew on crime/communion; and SANER is the only method that consistently reduces race–agency. Improvements are not single-axis: in some checkpoints, reducing gender–crime with Bias Prompts coincides with an increase in race–crime, indicating method-specific trade-offs. Effects depend on model size: smaller encoders achieve the most fair outcomes on gender, whereas reductions in racial skew are larger for ViT-L/14.
@@ 4.2 Encoder size
Increasing parameters from ViT-B/32 (63 M) to ViT-L/14 (428 M) while keeping the corpus fixed at 400 M pairs has opposite consequences for the two families (Figure 1). Within CLIP, the bigger encoder cuts gender skew by roughly one-third and leaves race skew statistically flat. Within OpenCLIP, the same scale-up amplifies both gender and race skew: gender skew rises by +0.29+0.29 on LAION-400M and by +0.18+0.18 on LAION-2B, while race skew climbs by +0.11+0.11 and +0.14+0.14, respectively. Parameter count, therefore, cannot be treated as an unconditional fairness lever; its effect depends on how those extra parameters interact with the training data.
Debiased models.
Model size modulates mitigation. In CLIP, Bias Prompts substantially reduce gender–crime for ViT-B/32 but become unstable for ViT-L/14, where they may exacerbate gender skew on PATA, even while reducing race–crime on FairFace. In OpenCLIP-400M, all three methods reduce gender skew more strongly for ViT-B/32, whereas reductions in racial skew are more pronounced for ViT-L/14. Aggregated across tasks, debiasing tends to reduce gender skew on smaller models and racial skew on larger models, suggesting increased risk of over-correction for gender-oriented prompts as model size increases.
@@ 4.3 Pre-training Scale
Debiased models.
Increasing LAION from 400M to 2B changes which strategy is most effective. For gender–crime, Prompt Array achieves the most fair outcomes on average (surpassing Bias Prompts at 400M); for race–crime, SANER becomes the most effective. Communion is broadly reduced by all strategies, while Agency remains brittle—on FairFace ViT-L/14, each method can overshoot after scaling. A plausible mechanism is that larger corpora improve lexical coverage (benefiting Prompt Array) and provide more neutral textual contexts (benefiting SANER), while additional data can also amplify racial bias that text-only adjustments cannot fully neutralize.
@@
4.4 Data Source
A direct comparison between CLIP and OpenCLIP at matched encoder size (B/32 or L/14) and corpus size (400 M) isolates the effect of swapping the proprietary crawl for LAION. OpenCLIP is systematically more gender-biased—by +0.07+0.07 on the small encoder and +0.18+0.18 on the large one—while the picture for race bias flips with model capacity. At 63 M parameters, OpenCLIP is less race-biased than CLIP; at 428 M parameters, it surpasses CLIP’s race skew (Figure 3). These complementary weaknesses suggest that the two corpora emphasise different social regularities and that ensemble or calibration methods may be needed to balance harms across demographic axes.
Debiased models.
Mitigation efficacy depends on training-data source. CLIP-tuned Bias Prompts reduce gender–agency on CLIP-B/32 but often amplify racial bias or gender–communion on OpenCLIP-400M. At 400M, OpenCLIP attains greater bias reduction with Prompt Array (race crime/communion) and with SANER (race–agency), whereas CLIP achieves its most fair outcomes with Bias Prompts. These differences likely reflect tokenization and embedding-geometry disparities and distinct visual pre-training statistics, which alter how text-space interventions propagate to image–text similarity.
@@ 4.5 Absolute harm frequencies
Debiased models.
Under mitigation, negative Communion is the most tractable: Prompt Array and SANER typically reduce its frequency more than Bias Prompts. Agency remains brittle—especially on ViT-L/14@FairFace—with occasional overshoot. For Crime/Animal, text-space debiasing is most effective on OpenCLIP, where Prompt Array and SANER yield the lowest mislabelling. Importantly, reductions on one axis (e.g., gender–Crime) can coincide with increases on another (e.g., race–Crime); we therefore track harm frequencies alongside skew to detect bias transfer and avoid amplifying total error mass.
Figure 4: Corpus-level prevalence of harmful top-1 predictions on 10,000 FairFace images. Each cluster shows the share of portraits tagged Criminal (yellow), negative Communion (orange), or negative Agency (red). Negative-communion stereotypes dominate—reaching 62% for CLIP-B/32@400M. Crime mislabelling tops out at 13% (CLIP-L/14@400M), while negative-agency labels peak at 36% (OpenCLIP-L/14@2B). Shorter bars indicate safer behaviour; whiskers give 95% bootstrap confidence intervals.
4.6 Debiasing strategies and debiased CLIP models
What works where. For CLIP, Bias Prompts are the only strategy that can reduce bias, primarily for ViT-B /32; for ViT-L/14, their effects are variable (e.g., large increases in gender–crime on PATA) despite reductions in some racial metrics on FairFace. For OpenCLIP-400M, Prompt Array is most effective at reducing racial skew on crime/communion, and SANER reduces racial–agency; Bias Prompts can exacerbate bias on FairFace. After scaling to 2B, Prompt Array becomes most effective for reducing gender–crime and SANER yields the largest reductions for race–crime.
Figure 5: CLIP vs. CLIP debiased with Bias Prompts. Panels: (a) Gender–Crime, (b) Gender–Communion, (c) Race–Crime, (d) Race–Communion. Bars show Bias Score (lower is better) across four checkpoints (B/32 $F$, L/14 $F$, B/32 PT, L/14 PT). Bias Prompts reduce skew in several settings but effects vary by axis and size, motivating validation at the deployment model size.
Figure 6: OpenCLIP-400M: original vs. three test-time debiasers. Panels: (a) Gender–Crime, (b) Gender–Communion, (c) Race–Crime, (d) Race–Communion. We compare the original model with Bias Prompts, SANER, and Prompt Array. Method efficacy is source- and size-dependent; per-axis gains should be monitored for displacement.
Figure 7: OpenCLIP-2B: original vs. Bias Prompts, SANER, and Prompt Array. Panels: (a) Gender–Crime, (b) Gender–Communion, (c) Race–Crime, (d) Race–Communion. Scaling to 2B alters which intervention helps most; improvements on one axis can coincide with regressions on another.
Figure 8: Family comparison under Bias Prompts. Panels: (a) Gender–Crime, (b) Gender–Communion, (c) Race–Crime, (d) Race–Communion. Side-by-side Bias Scores for CLIP and OpenCLIP in original and debiased (Bias Prompts) conditions highlight family-specific responses to the same intervention.
Encoder scale, corpus scale, and corpus source each leave a distinct fingerprint on bias; debiasing partially offsets—but does not erase—these patterns. Larger proprietary-data CLIP models reduce gender skew while leaving race roughly unchanged, whereas larger LAION-trained OpenCLIP models amplify both gender and racial bias; increasing LAION size further exacerbates bias, and swapping proprietary data for LAION trades gender fairness for higher racial skew at matched budgets. Against this backdrop, Bias Prompts are most effective for reducing gender skew in CLIP-B/32, while Prompt Array and SANER more reliably reduce racial skew in OpenCLIP; scaling from 400M to 2B alters which strategy is most fair. Hence, mitigation must be selected per metric and model family and validated at the deployment model size, lest an intervention reduce bias on one axis while transferring or amplifying harm on another.
@@ 5 Discussion
Mitigation effects mirror the family-specific scaling asymmetry. Prompt-level interventions tend to reduce gender skew on smaller models but become less stable as model size increases; by contrast, reductions on racial axes are more likely to materialise on larger models. These patterns suggest that text-space adjustments interact non-linearly with representational capacity and should be validated at the deployment model size, where over-correction can exacerbate bias on untargeted axes.
@@
A five-fold jump from LAION-400M to LAION-2B delivers far greater coverage of visual concepts, yet every measured bias metric moves in the wrong direction for OpenCLIP. The larger crawl simply repeats the demographic profile of its smaller sibling, giving majority groups five times as many updates and elevating their linguistic footprints in the contrastive objective. Size alone, therefore, fails to dilute bias; without careful rebalancing, scale can ossify it.
Increasing training-data scale reconfigures which strategies are most effective. For larger LAION crawls, array-based prompting often attains the most fair outcomes on gender–Crime, while neutralisation-based methods are more effective for race–Crime. Communion is comparatively tractable across methods, whereas Agency remains brittle. Expanded corpora improve lexical coverage and neutral contexts for text-space mitigation, but can simultaneously amplify racial bias patterns that lightweight interventions do not fully neutralise.
Source outweighs raw scale.
Even at identical encoder size and corpus size, swapping the proprietary WebImageText crawl for LAION flips the fairness profile: OpenCLIP becomes more gender-biased and, once the model is large enough, more race-biased as well. These shifts highlight that what matters is not merely how many image–text pairs a model sees but which pairs—and whose voices—populate the crawl. Corpus source thus emerges as a lever at least as powerful as parameter or sample count.
Mitigation is source-dependent. Strategies tuned on CLIP typically reduce gender skew for CLIP but can amplify racial bias or Communion skew for OpenCLIP at matched budgets. OpenCLIP more reliably benefits from array-based and neutralisation-based approaches on racial metrics. Tokenisation, embedding geometry, and visual pre-training statistics differ across corpora, so prompts and neutralisers should be re-tuned rather than transferred verbatim.
@@
Improvements are not uniformly shared across axes: reducing gender–Crime can coincide with increased race–Crime. To avoid shifting or amplifying harm, debiasing should be assessed with both disparity metrics and harm-frequency measures, and coupled with simple calibration or representation-level regularisation when reversals are observed.
@@
To turn these diagnostics into durable mitigation, adopt principled, scale-stable debiasing objectives and corpus processes that are (i) source-robust across crawls, (ii) size-stable as parameters and data grow, (iii) axis-coherent—reducing gender and racial skew without displacement—and (iv) preserve explicitly specified attributes. In practice, favour corpus-aware re-tuning and sample-efficient calibration over ad hoc prompt tweaks; for risk-sensitive deployments, pair light regularisers with continuous monitoring of all six axes at the target model size.
@@ 6 Conclusion
By disentangling encoder size, pre-training scale, and corpus source while holding the contrastive loss fixed, this paper shows that each lever—and their interactions—imprints a distinct bias profile on CLIP-style vision–language models. Increasing model size moves proprietary-data CLIP toward gender parity while leaving race roughly unchanged, whereas the same increase in OpenCLIP amplifies both gender and racial skew. Expanding LAION from 400M to 2B further increases bias, and at matched size and scale, replacing proprietary data with LAION trades improved gender fairness for heightened racial skew. Fairness is therefore not an automatic by-product of scaling; it emerges from how additional capacity couples with the statistical structure of the data.
We also evaluate three post-hoc, test-time debiasing strategies. While debiasing reduces harm, its effectiveness is source- and size-dependent: strategies that reduce gender skew in CLIP do not necessarily reduce racial skew in OpenCLIP, and changes in data scale can invert effectiveness. These observations suggest that selecting a method per task and model is, at best, a stopgap. What is needed are principled, scale-stable debiasing objectives and data processes that (i) are source-robust across corpora, (ii) remain size-stable as models and datasets grow, (iii) deliver axis-coherent improvements (reducing gender and racial skew without transferring harm), and (iv) preserve attribute information when explicitly specified.
Looking forward, we argue for fairness as a first-class design constraint: corpus-aware pre-training with demographic rebalancing and counterfactual augmentation; contrastive objectives regularised for group-invariant similarity and distributional robustness; lightweight, text–image neutralisers with guarantees on harm-transfer; and multi-objective optimisation that explicitly trades off accuracy, disparity, and harm frequency under a single training budget. Rather than choosing prompts to mask symptoms, future VLMs should be trained and evaluated under protocols where fairness is built in and scale-stable. Our evaluation suite, code, and controls are released to make such development routine and reproducible, and to guide the construction of open-vocabulary systems that serve diverse users equitably.
Appendix A Neutral-Image Control
To verify that our bias metrics do not spuriously signal structure on demographically–neutral inputs, we generated 19 feature-neutral portraits with DALLE and Gemini (e.g. “average composite face, facial features blurred, grey background”). Figure 9 shows the full set.
Each portrait was embedded by six checkpoints (CLIP B/32, L/14 and OpenCLIP B/32, L/14 trained on LAION-400M and LAION-2B). For every model we report:
  1. (1)
    Intra-set cosine $(\mu,\sigma)$ – mean and standard deviation of pair-wise cosine similarity between the 19 portraits (summary sheet); and
  2. (2)
    Directional bias $\Delta=\mu_{\text{pos}}-\mu_{\text{neg}}$ for the communal (COMM), agentic (AGEN) and crime (CRIME) prompt families (directional_bias sheet).
Table 5: Neutral-image sanity check (all numbers $\times 10^{-2}$). Directional bias collapses to $\lvert\Delta\rvert\leq 0.02$ for COMM and AGEN and $\leq 0.05$ for CRIME, an order of magnitude below the skews observed on FairFace and PATA.
Model Intra-set cosine Directional bias $\downarrow$
$\mu$ $\sigma$ COMM AGEN CRIME
CLIP ViT-B/32 73.2 9.4 -0.01 -0.34 -0.21
CLIP ViT-L/14 68.3 10.5 -0.54 -1.12 +0.27
OpenCLIP ViT-B/32400M 57.1 13.9 -0.30 -2.17 -1.70
OpenCLIP ViT-B/322B 57.5 12.7 -0.04 -1.61 -2.36
OpenCLIP ViT-L/14400M 51.9 13.4 +0.53 -1.49 -5.14
OpenCLIP ViT-L/142B 52.0 12.9 -0.15 -0.99 -2.56
Interpretation.
High cosine means (>0.5>0.5) show the portraits cluster tightly, while directional bias remains near zero: $\lvert\Delta_{\text{COMM/AGEN}}\rvert\!\leq\!0.02$ and $\lvert\Delta_{\text{CRIME}}\rvert\!\leq\!0.05$. These values establish the intrinsic floor of our metric.
Figure 9: Nineteen feature-neutral portraits used for the control experiment. No visible race or gender cues are present, yet the set spans diverse lighting and artistic styles.
Appendix B Calibration of Directional-Bias Magnitude
Goal.
We calibrate directional bias magnitudes to practical risk by mapping the absolute cosine difference $|\Delta|$ to the probability that the model outputs a harmful label: the negative-pole prompt for Comm/Agen, or an animal/criminal term for Crime.
Method.
For each PATA face, we bin $|\Delta|$ into seven ranges333[0.0, 0.1), [0.1, 0.2), …, [0.6, $\infty$). and estimate the empirical For each PATA face, we bin $|\Delta|$ into seven ranges444[0.0, 0.1), [0.1, 0.2), …, [0.6, $\infty$). and estimate the empirical
[…]
Reading the curve.
A flat segment near $p_{\mathrm{harm}}\!\approx\!0.1$ indicates chance-level behavior. The elbow is the first bin with $p_{\mathrm{harm}}\!\geq\!0.5$; the plateau indicates saturation. Figure 10 shows a typical shape.
Figure 10: Calibration curve for CLIP ViT-B/32 on Comm. The elbow occurs in the 0.2–0.3 bin and the curve saturates at $p_{\mathrm{harm}}=0.91$ for $|\Delta|\geq 0.5$.
Aggregated results.
Table 6 summarizes the elbow bin and the maximal $p_{\mathrm{harm}}$ for each checkpoint and task.
Table 6: Directional-bias calibration summary. Lower elbows and higher maxima indicate greater risk. The elbow is the smallest $|\Delta|$ bin with $p_{\mathrm{harm}}\geq 0.5$; “—” means no bin reaches 0.5.
Comm Agen Crime
Model elbow max $p_{\mathrm{harm}}$ elbow max $p_{\mathrm{harm}}$ elbow max $p_{\mathrm{harm}}$
CLIP ViT-B/32 0.2–0.3 0.91 0.3–0.4 0.83 0.3–0.4 0.94
CLIP ViT-L/14 0.15 0.4–0.5 0.64 0.2–0.3 0.72
OpenCLIP B/32400M 0.1–0.2 0.96 0.2–0.3 0.97 0.2–0.3 0.99
OpenCLIP B/322B 0.2–0.3 0.93 0.2–0.3 0.95 0.2–0.3 0.98
OpenCLIP L/14400M 0.1–0.2 0.98 0.1–0.2 0.99 0.1–0.2 1.00
OpenCLIP L/142B 0.1–0.2 0.95 0.2–0.3 0.97 0.2–0.3 0.99
Interpretation.
(i) Bias magnitudes below $0.1$ are effectively benign: $|\Delta|<0.1$ shifts $p_{\mathrm{harm}}$ by <2<\!2 pp, matching neutral-image controls. (ii) For CLIP ViT-B/32, risk crosses $50\%$ around $|\Delta|\!\approx\!0.25$ and reaches $\geq\!90\%$ beyond $0.5$, marking moderate vs. severe bias thresholds. (iii) OpenCLIP curves are steeper: elbows often appear in the 0.1–0.2 bin and $p_{\mathrm{harm}}\!\geq\!95\%$ by $|\Delta|=0.3$, consistent with the larger max-skew values in the main text. (iv) Crime saturates fastest, indicating that non-human or criminal mislabellings are particularly sensitive to directional bias.
These calibration curves turn cosine magnitudes into actionable thresholds: a seemingly small $|\Delta|=0.3$ already implies a four-in-five chance of a harmful prediction for several checkpoints, making directional bias a practically meaningful metric.
Appendix C Robustness to Prompt-Template Variation
Motivation.
Bias estimates could in principle, depend on the exact wording of the zero-shot prompt. We therefore repeat the PATA audit under three template families: 1. orig: “a photo of a {label} (main paper), 2. image_of: “an image of a {label}, 3. portrait: “portrait of a {label}.
Protocol.
For every model–task pair, we re-encode all 3,948 faces and record the directional-bias value $\Delta$. Table 7 reports the mean ( $\mu$) and standard deviation ( $\sigma$) across the dataset.
Table 7: Template sensitivity of the bias metric. $\Delta\mu$ = max absolute difference in means across the three templates. All deltas are $\leq 0.02$ cosine units, far below the effect sizes discussed in §4.
COMM AGEN CRIME
Model $\mu$ $\Delta\mu$ $\mu$ $\Delta\mu$ $\mu$ $\Delta\mu$
CLIP ViT-B/32 0.042 0.008 0.039 0.009 0.065 0.010
CLIP ViT-L/14 0.028 0.006 0.027 0.007 0.048 0.011
OpenCLIP B/32400M 0.071 0.011 0.078 0.010 0.095 0.013
OpenCLIP B/322B 0.066 0.010 0.074 0.009 0.090 0.011
OpenCLIP L/14400M 0.083 0.012 0.088 0.013 0.162 0.017
OpenCLIP L/142B 0.079 0.011 0.081 0.012 0.149 0.016
Findings.
Changing the prompt wording alters the absolute bias by at most $0.02$ cosine units ( $\approx$5 % of the effect sizes between models). Rank order among checkpoints is unchanged, and the neutral-image control (§A) retains $|\Delta|\!<\!0.05$ under all templates. Hence, the bias patterns we report are template-stable and not an artefact of a particular phrase.
Appendix D Experimental Setup
@@
D.1 Attribute Prediction Bias